# Ruishu Liu, Yulin Cao,  Xiaojie Wang

### v1

@@ Abstract
The present work is devoted to strong approximations of a generalized Aït-Sahalia model arising from mathematical finance. The numerical study of the considered model faces essential difficulties caused by a drift that blows up at the origin, highly nonlinear drift and diffusion coefficients and positivity-preserving requirement. In this paper, a novel explicit Euler-type scheme is proposed, which is easily implementable and able to preserve positivity of the original model unconditionally, i.e., for any time step-size $h>0$. A mean-square convergence rate of order $0.5$ is also obtained for the proposed scheme in both non-critical and general critical cases. Our work is motivated by the need to justify the multi-level Monte Carlo (MLMC) simulations for the underlying model, where the rate of mean-square convergence is required and the preservation of positivity is desirable particularly for large discretization time steps. To the best of our knowledge, this is the first paper to propose an unconditionally positivity preserving explicit scheme with order $1/2$ of mean-square convergence for the model. Numerical experiments are finally provided to confirm the theoretical findings.
Keywords: Aït-Sahalia model; Explicit scheme; Unconditionally positivity preserving; Strong convergence rates.
@@ 1 Introduction
The pricing of financial assets has been an important issue in mathematical finance over the past few decades, where stochastic differential equations (SDEs) are prominent models frequently used to capture the behaviors of financial systems. Notable examples include the Black-Scholes model [2], the Merton model [14] and the Cox-Ingersoll-Ross (CIR) model [5] in the modern financial theory. The economist and mathematician Aït-Sahalia conducted an empirical study in [1] and investigated several continuous-time models for interest rates. He tested the models by comparing the parameter density implied by the parameter models, and found that all the existing univariate linear drift models fail to explain the euro dollar well. As a consequence of this work, a nonlinear stochastic interest model which captures well of the dynamics of the spot rate was proposed and named as the Aït-Sahalia model, which is now widely used in finance and economics, and described by the following Itô-type SDE:
@@
Here $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\rho>1$ and $\{W_{t}\}_{t\in[0,T]}$ stands for a standard Brownian motion defined on a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Aït-Sahalia [1] proved that the model (1.1) admits a unique global positive solution using the Feller explosion test. In 2009, Cheng [4] derived the finite moment bounds for the analytic solutions and proved that the Euler–Maruyama approximate solutions are convergent in probability. Later in 2011, Szpruch et al. [18] extended the model (1.1) to a more general version given by
$\mathrm{d}X_{t}=$ $(c_{-1}X_{t}^{-1}-c_{0}+c_{1}X_{t}-c_{2}X_{t}^{\kappa})\,\text{d}t+c_{3}X_{t}^{\rho}\text{d}W_{t},\quad X_{0}=x_{0}>0,\>\kappa,\rho>1,$ (1.2)
and showed that the generalized model admits a unique global positive solution with uniformly in time bounded $p$-th moments ($p\in\mathbb{R}$) in the non-critical case $\kappa+1>2\rho$. Moreover, they applied the classical backward Euler-Maruyama (BEM) scheme to approximate the model and proved the positivity-preserving approximations to be strongly convergent under the condition that $\kappa+1>2\rho$, but with no convergence rates revealed. Recently, the authors of [20] successfully recovered the mean-square convergence rate of order $1/2$ for the stochastic theta methods (STMs) with $\theta\in[\tfrac{1}{2},1]$ applied to the model (1.2) under the condition that $\kappa+1\geq 2\rho$, covering the critical case $\kappa+1=2\rho$. Also, we mention that Neuenkirch and Szpruch [15] proposed a kind of Lamperti-backward Euler method for scalar SDEs defined in a domain, including the Ait-Sahalia-type model (1.2), with a strong convergence rate of order one obtained. However, the computational costs of solving implicit algebraic equations rise as the parameter $\kappa$ increases. More recently, a truncated explicit Euler method was proposed in [7] to approximate (1.2), but with no convergence rate revealed. When we finish the draft of the present work, we are also aware of two closely related publications [6, 13]. The authors of [6] recovered a mean-square convergence rate of order nearly $\tfrac{1}{4}$ for the truncated Euler method only in the non-critical case $\kappa+1>2\rho$. The work [13] proposed an exponential tamed method based on the Lamperti transform and splitting, which is strongly convergent with order one, but not strictly positivity preserving.
In the present article we aim to design novel explicit positivity preserving schemes for approximations of the Aït-Sahalia model (1.2) with mean-square convergence rate of order $\tfrac{1}{2}$. More precisely, on a uniform mesh with a uniform step-size $h=\tfrac{T}{N},\,N\in\mathbb{N}$ we propose the following time-stepping scheme:
@@
where $f_{h}$ and $g_{h}$ are some explicit modifications of $f(x):=-c_{2}x^{\kappa}$ and $g(x):=c_{3}x^{\rho}$ in the SDE model, respectively (see Section 3 for details). Clearly, the implicitness only appears in the term $c_{-1}Y_{n+1}^{-1}h$, which is the key element to preserve the positivity of the original model (see Lemma 3.1). On the contrary, the remaining terms are explicitly handled. It is not difficult to see that, given the former step $Y_{n}$, the current step $Y_{n+1}$ can be explicitly known, as the unique positive root of a quadratic equation (see (3.5) below for the explicit solution). In this way, the computational costs are significantly reduced compared with the implicit schemes (BEM or STMs) in existing publications [18, 20], where implicit algebraic equations are required to be numerically solved for every step. Furthermore, the proposed explicit Euler-type scheme is positivity-preserving for any time step-size $h>0$, which is particularly desirable for the multi-level Monte Carlo (MLMC) simulations [9] with large discretization time steps.
To justify the MLMC simulations, the mean-square convergence rate of the approximations is also necessary [9, Theorem 3.1]. Recall that the analysis of convergence rates in [20] for the implicit schemes does not rely on a priori moment estimates of numerical approximations. Nevertheless, due to the presence of the explicit treatment in the drift term, the analysis of convergence rates for the proposed semi-implicit schemes relies on finite moment bounds of numerical approximations:
@@
Deriving such $p$-th moment integrability of the numerical approximations is, however, not an easy task and some non-standard arguments are exploited (see the proof of Lemma 3.6), to overcome essential difficulties caused by a drift that blows up at the origin, highly nonlinear diffusion coefficients and a mixture of implicitness and explicitness in the drift part of the scheme. Equipped with the moment bound (1.4), one is able to identify the mean-square convergence rate of order $0.5$ for the proposed scheme (3.1) applied to the model (1.2) (see Theorem 4.1):
@@
in both non-critical ($\kappa+1>2\rho$) and general critical ($\kappa+1=2\rho$) cases, where the constant $C$ is independent of the time step-size $h$. To the best of our knowledge, this is the first paper to devise an unconditionally positivity preserving explicit scheme with order $\tfrac{1}{2}$ of mean-square convergence for the general Aït-Sahalia model.
The rest of this paper is organized as follows. In the next section, basic notations and properties of the original model are presented. In Section 3 the newly proposed scheme is introduced and proved to be positivity preserving and moment bounded. The mean-square error bounds for the proposed scheme is carefully analyzed in Section 4, with order one-half of convergence revealed. Numerical experiments are shown in Section 5 to verify our theoretical results and a short conclusion is made in Section 6.
@@ 2 The generalized Aït-Sahalia model
$\|\xi\|_{\mathcal{L}^{p}(\Omega;\mathbb{R})}:=\big{(}\mathbb{E}\big{[}|\xi|^{p}\big{]}\big{)}^{1/p}<\infty.$ (2.1)
In this work we focus on the generalized Aït-Sahalia model taking the form as follows:
@@
where $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\kappa$, $\rho>1$ and $\kappa+1\geq 2\rho$. For the particular case $\kappa=2$, the underlying model (2.2) reduces into (1.1), originally proposed by Aït-Sahalia [1]. The well-posedness of the generalized model (2.2) has been established in the following theorem quoted from [18, Theorem 2.1].
@@
Furthermore, we recall the moment integrability properties of the solution to (2.2) as follows (see [18, Lemma 2.1] and [20, Lemma 4.6] for details).
@@
[…]
For the general critical case $\kappa+1=2\rho$, we have slightly different assertions, as stated in the following lemma (cf. [20, Lemma 4.6]).
@@
Let conditions in Theorem 2.1 be satisfied with $\kappa+1=2\rho$. Then for any $12\leq p_{1}\leq\tfrac{2c_{2}}{c_{3}^{2}}+$ 1 and for any $p_{2}\geq 2$ we have
[…]
@@
[20, Lemma 4.4] Let conditions in Theorem 2.1 be satisfied with $\kappa+1>2\rho$. Then it holds that, for any $p\geq 1$ and $t,s\in[0,T]$,
@@
[20, Lemma 4.7] Let conditions in Theorem 2.1 be satisfied with $\kappa+1=2\rho$. Then for any $t,s\in[0,T]$ it holds that
@@
Here and below, for notation simplicity the letter $C$ is slightly abused to denote a generic positive constant, which is independent of the time step-size and may vary for each appearance.
As a straightforward consequence of the above lemmas, one can use the Hölder inequality to show the following estimates (consult [20, (4.21), (4.24)]).
@@
then for any $t,s\in[0,T]$ it holds that
@@ 3 The explicit positivity-preserving Euler-type scheme
In this section we aim to propose and analyze new strong approximation schemes for the generalized Aït-Sahalia model (2.2). For $T\in(0,+\infty)$ and $N\in\mathbb{N}$, we construct a uniform mesh $\{t_{n}=nh\}_{n=0}^{N}$ on the time interval $[0,T]$ with a uniform step-size $h=\tfrac{T}{N}$. Based on the uniform mesh, we propose an explicit Euler-type method for (2.2) as follows:
@@
Let $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\kappa$, $\rho>1$. Given any initial value $Y_{0}=X_{0}>0$, the scheme (3.1) admits unique, positive numerical solutions $\{Y_{n}\}_{n\in\mathbb{N}}$ for any stepsize $h=\tfrac{T}{N}>0$. In other words, the scheme (3.1) is positivity preserving unconditionally.
Proof: For $x>0$ we introduce the function
@@
Let $\kappa+1\geq 2\rho$. There exist a positive constant $L(\gamma)$, independent of the stepsize $h>0$, such that
@@
where $\gamma>0$ is assumed to be arbitrarily large for the case $\kappa+1>2\rho$ and $\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ for the critical case $\kappa+1=2\rho$.
@@
There exists some positive constant $m_{1},m_{2}\leq 2\gamma+1$ such that
@@
where $\alpha$ is a positive constant satisfying $\alpha\geq\tfrac{1}{2}$. Here the choice of $f_{h}$ and $g_{h}$ are nothing but a kind of taming strategy, which was firstly proposed in [10] and later extensively studied by many authors, see, e.g., [19, 3, 16, 17, 12, 13]. To validate Assumption 3.2, noticing that $2\kappa\alpha\geq\kappa>\rho$, simple calculations infer that for any $x\in(0,+\infty)$
@@
Thus we have
[…]
@@
$\big{|}g(x)-g_{h}(x)\big{|}$ $=\tfrac{c_{3}h^{\alpha}x^{(2\alpha+1)\rho}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq c_{3}h^{\alpha}x^{(2\alpha+1)\rho}.$ (3.15)
Let $m_{1}=2(2\alpha+1)\kappa$ and $m_{2}=2(2\alpha+1)\rho$. Since $\alpha\geq\tfrac{1}{2}$, (3.8) holds true on an additional condition $12(2\alpha+1)\kappa\leq 2\gamma+$ 1.
@@
Let Assumptions 3.2, 3.3 hold. Then for any $12\leq p\leq 2\gamma+$ 1 and for any step-size $h>0$ it holds that
@@
Next we introduce a time-continuous version of $Z_{n+1}$ in each time interval $[t_{n},t_{n+1}]$ as follows:
@@
Note that $\widetilde{Z}_{t_{n}}^{(n)}=Y_{n}\neq Z_{n}=\widetilde{Z}_{t_{n}}^{(n-1)}$. Owing to (3.6), a rough moment bound can be derived for $\widetilde{Z}_{t}^{(n)}$. Indeed, by the Jensen inequality, for any $p\geq 1$ we have
@@
where $\varepsilon\geq 2c_{-1}h>0$ is a positive constant specified later and $\hat{C}$$,\tilde{C}_{\varepsilon}$ are positive constants independent of $h$. Thus
$\mathbb{E}\bigg{[}\Big{(}\varepsilon+|Y_{n+1}|^{2}+c_{-1}^{2}h^{2}|Y_{n+1}|^{-2}-2c_{-1}h\Big{)}^{p}\bigg{]}=\mathbb{E}\bigg{[}\Big{(}\varepsilon+\big{|}Y_{n+1}-c_{-1}hY_{n+1}^{-1}\big{|}^{2}\Big{)}^{p}\bigg{]}\leq\hat{C}\mathbb{E}\Big{[}\big{|}Y_{n}\big{|}^{2p}\Big{]}+\tilde{C}_{\varepsilon}.$ (3.21)
By noting $\varepsilon\geq 2c_{-1}h$, it holds that
@@
By iteration we get
$\sup_{0\leq n\leq N}\mathbb{E}\big{[}\big{|}Y_{n}\big{|}^{2p}\big{]}<\infty,$ (3.23)
which, together with (3.20), infers that
@@
We emphasize that the above rough moment bound, depending on $N$, is not proved to be uniformly bounded with respect to $N$. Nevertheless, it suffices to allow the use of the Gronwall lemma later to arrive at uniform moment bounds. To do this, we let $\varepsilon\geq 1+2c_{-1}h$ in the following and denote
$V_{\varepsilon,q}(x):=\big{(}\varepsilon+|x|^{2}\big{)}^{q},\quad q,x\in\mathbb{R}.$ (3.25)
It can be easily checked that for any $q,q^{\prime}\in\mathbb{R}$,
@@
$V_{\varepsilon,q}^{\prime\prime}(x)$ $=2q\big{(}\varepsilon+|x|^{2}\big{)}^{q-2}\big{[}\varepsilon+(2q-1)|x|^{2}\big{]}$ $\leq 2q\cdot\max{\{1,2q-1\}}\cdot V_{\varepsilon,q-1}(x).$
@@
Equipped with these properties, we can start the estimation of $\widetilde{Z}_{t}^{(n)}$. The Itô formula together with (3.28) and (3.29) yields
@@
where we also used (3.19) in the last step. In view of (3.7), for any $q\leq\gamma+\tfrac{1}{2}$, one deduces that
@@
Observing that
$\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s}\Bigg{]}=0,$ (3.35)
taking expectations on both sides of (3.34) reveals that

### v2

@@ Abstract
The present work is devoted to strong approximations of a generalized Aït-Sahalia model arising from mathematical finance. The numerical study of the considered model faces essential difficulties caused by a drift that blows up at the origin, highly nonlinear drift and diffusion coefficients and positivity-preserving requirement. In this paper, a novel explicit Euler-type scheme is proposed, which is easily implementable and able to preserve positivity of the original model unconditionally, i.e., for any time step-size $h>0$. A mean-square convergence rate of order $0.5$ is also obtained for the proposed scheme in both non-critical and general critical cases. Our work is motivated by the need to justify the multi-level Monte Carlo (MLMC) simulations for the underlying model, where the rate of mean-square convergence is required and the preservation of positivity is desirable particularly for large discretization time steps. Numerical experiments are finally provided to confirm the theoretical findings.
Keywords: Aït-Sahalia model; Explicit scheme; Unconditionally positivity preserving; Strong convergence rate.
@@ 1 Introduction
The pricing of financial assets has been an important issue in mathematical finance over the past few decades, where stochastic differential equations (SDEs) are prominent models frequently used to capture the behaviors of financial systems. Notable examples include the Black-Scholes model [2], the Merton model [16], and the Cox-Ingersoll-Ross (CIR) model [5] in the modern financial theory. The economist and mathematician Aït-Sahalia conducted an empirical study in [1] and investigated several continuous-time models for interest rates. He tested the models by comparing the parameter density implied by the parameter models and found that all the existing univariate linear drift models fail to explain the euro dollar well. As a consequence of this work, a nonlinear stochastic interest model which captures well of the dynamics of the spot rate was proposed and named as the Aït-Sahalia model, which is now widely used in finance and economics, and described by the following Itô-type SDE:
@@
Here, $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\rho>1$ and $\{W_{t}\}_{t\in[0,T]}$ stands for a standard Brownian motion defined on a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Aït-Sahalia [1] proved that the model (1.1) admits a unique global positive solution using the Feller explosion test. In 2009, Cheng [4] derived the finite moment bounds for the analytic solutions and proved that the Euler–Maruyama approximate solutions are convergent in probability. Later in 2011, Szpruch et al. [20] extended the model (1.1) to a more general version given by
$\mathrm{d}X_{t}=$ $(c_{-1}X_{t}^{-1}-c_{0}+c_{1}X_{t}-c_{2}X_{t}^{\kappa})\,\text{d}t+c_{3}X_{t}^{\rho}\text{d}W_{t},\quad X_{0}=x_{0}>0,\>\kappa>1,$ (1.2)
and showed that the generalized model admits a unique global positive solution with uniformly in time bounded $p$-th moments ($p\in\mathbb{R}$) in the non-critical case $\kappa+1>2\rho$. Moreover, they applied the classical backward Euler-Maruyama (BEM) scheme to approximate the model and proved the positivity-preserving approximations to be strongly convergent under the condition that $\kappa+1>2\rho$, but with no convergence rates revealed. Recently, the authors of [22] successfully recovered the mean-square convergence rate of order $1/2$ for the stochastic theta methods (STMs) with $\theta\in[\tfrac{1}{2},1]$ applied to the model (1.2) under the condition that $\kappa+1\geq 2\rho$, covering the critical case $\kappa+1=2\rho$. Also, we mention that Neuenkirch and Szpruch [17] proposed a kind of Lamperti-backward Euler method for scalar SDEs defined in a domain, including the Ait-Sahalia-type model (1.2), with a strong convergence rate of order one obtained. However, the computational costs of solving implicit algebraic equations rise as the parameter $\kappa$ increases. More recently, a truncated explicit Euler method was proposed in [7] to approximate (1.2), but with no convergence rate revealed. When we finish the draft of the present work, we are also aware of two closely related publications [6, 14]. The authors of [6] recovered a mean-square convergence rate of order nearly $\tfrac{1}{4}$ for the truncated Euler method only in the non-critical case $\kappa+1>2\rho$. The work [14] proposed an exponential tamed method based on the Lamperti transform and splitting, which is strongly convergent with order one, but not strictly positivity preserving. Finally, it is worthwhile to mention an interesting paper [10], where, by combining a Lamperti-type transformation with a semi-implicit approach, the authors proposed a positivity preserving scheme for the model in the particular critical case $\kappa=2,\rho=\tfrac{3}{2}$. For the particular critical case, the proposed scheme can be explicitly solved, by finding a positive root of a quadratic equation. Moreover, it was shown to have order one strong convergence (see [10] for more details).
In the present article, we aim to design novel explicit positivity preserving schemes for approximations of the Aït-Sahalia model (1.2) with a mean-square convergence rate of order $\tfrac{1}{2}$. More precisely, on a uniform mesh with a uniform step-size $h=\tfrac{T}{N},\,N\in\mathbb{N}$ we propose the following time-stepping scheme:
@@
where $f_{h}$ and $g_{h}$ are some explicit modifications of $f(x):=-c_{2}x^{\kappa}$ and $g(x):=c_{3}x^{\rho}$ in the SDE model, respectively (see Sect. 3 for details). Clearly, the implicitness only appears in the term $c_{-1}Y_{n+1}^{-1}h$, which is the key element to preserve the positivity of the original model (see Lemma 3.1). On the contrary, the remaining terms are explicitly handled. It is not difficult to see that, given the former step $Y_{n}$, the current step $Y_{n+1}$ can be explicitly known, as the unique positive root of a quadratic equation (see (3.5) below for the explicit solution). In this way, the computational costs are significantly reduced compared with the implicit schemes (BEM or STMs) in existing publications [20, 22], where implicit algebraic equations are required to be numerically solved for every step. Furthermore, the proposed explicit Euler-type scheme is positivity-preserving for any time step-size $h>0$, which is particularly desirable for the multi-level Monte Carlo (MLMC) simulations [9] with large discretization time steps.
To justify the MLMC simulations, the mean-square convergence rate of the approximations is also necessary [9, Theorem 3.1]. Recall that the analysis of convergence rates in [22] for the implicit schemes does not rely on a priori moment estimates of numerical approximations. Nevertheless, due to the presence of the explicit treatment in the drift term, the analysis of convergence rates for the proposed semi-implicit schemes relies on a finite moment bound of numerical approximations:
@@
Deriving such $p$-th moment integrability of the numerical approximations is, however, not an easy task and some non-standard arguments are exploited (see the proof of Lemma 3.6) to overcome essential difficulties caused by a drift that blows up at the origin, highly nonlinear diffusion coefficients and a mixture of implicitness and explicitness in the drift part of the scheme. Equipped with the moment bound (1.4), one is able to identify the mean-square convergence rate of order $0.5$ for the proposed scheme (3.1) applied to the model (1.2) (see Theorem 4.1):
@@
in both non-critical ($\kappa+1>2\rho$) and general critical ($\kappa+1=2\rho$) cases, where the constant $C$ is independent of the time step-size $h$. To the best of our knowledge, this is the first paper to devise an unconditionally positivity preserving explicit scheme with order $\tfrac{1}{2}$ of mean-square convergence for the non-critical case of the Aït-Sahalia model (1.2).
The rest of this paper is organized as follows. In the next section, basic notations and properties of the original model are presented. In Sect. 3 the newly proposed scheme is introduced and proved to be positivity preserving and moment bounded. The mean-square error bounds for the proposed scheme is carefully analyzed in Sect. 4, with order one-half of convergence revealed. Numerical experiments are shown in Sect. 5 to verify our theoretical results, and a short conclusion is made in Sect. 6.
@@ 2 The generalized Aït-Sahalia model
$\|\xi\|_{L^{p}(\Omega;\mathbb{R})}:=\big{(}\mathbb{E}\big{[}|\xi|^{p}\big{]}\big{)}^{1/p}<\infty.$ (2.1)
In this work, we focus on the generalized Aït-Sahalia model taking the form as follows:
@@
where $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\kappa$, $\rho>1$ and $\kappa+1\geq 2\rho$. For the particular case $\kappa=2$, the underlying model (2.2) reduces into (1.1), originally proposed by Aït-Sahalia [1]. The well-posedness of the generalized model (2.2) has been established in the following theorem quoted from [20, Theorem 2.1].
@@
Furthermore, we recall the moment integrability properties of the solution to (2.2) as follows (see [20, Lemma 2.1] and [22, Lemma 4.6] for details).
@@
$\sup_{0\leq t<\infty}\mathbb{E}\big{[}|X_{t}|^{p}\big{]}<\infty\quad\text{and}\quad\sup_{0\leq t<\infty}\mathbb{E}\big{[}\big{|}\tfrac{1}{X_{t}}\big{|}^{p}\big{]}<\infty.$ (2.3)
For the general critical case $\kappa+1=2\rho$, we have slightly different assertions, as stated in the following lemma (cf. [22, Lemma 4.6]).
@@
Let conditions in Theorem 2.1 be satisfied with $\kappa+1=2\rho$. Then for any $12\leq p_{1}\leq\tfrac{2c_{2}}{c_{3}^{2}}+$ 1 and for any $p_{2}\geq 2$, we have
$\sup_{0\leq t<\infty}\mathbb{E}\big{[}|X_{t}|^{p_{1}}\big{]}<\infty\quad\text{and}\quad\sup_{0\leq t<\infty}\mathbb{E}\big{[}\big{|}\tfrac{1}{X_{t}}\big{|}^{p_{2}}\big{]}<\infty.$ (2.4)
@@
[22, Lemma 4.4] Let conditions in Theorem 2.1 be satisfied with $\kappa+1>2\rho$. Then it holds that, for any $p\geq 1$ and $t,s\in[0,T]$,
@@
[22, Lemma 4.7] Let conditions in Theorem 2.1 be satisfied with $\kappa+1=2\rho$. Then for any $t,s\in[0,T]$, it holds that
@@
Here and below, for notation simplicity, the letter $C$ is slightly abused to denote a generic positive constant, which is independent of the time step-size and may vary for each appearance.
As a straightforward consequence of the above lemmas, one can use the Hölder inequality to show the following estimates (consult [22, (4.21), (4.24)]).
@@
then for any $t,s\in[0,T]$, it holds that
@@ 3 The explicit positivity-preserving Euler-type scheme
In this section, we aim to propose and analyze new strong approximation schemes for the generalized Aït-Sahalia model (2.2). For $T\in(0,+\infty)$ and $N\in\mathbb{N}$, we construct a uniform mesh $\{t_{n}=nh\}_{n=0}^{N}$ on the time interval $[0,T]$ with a uniform step-size $h=\tfrac{T}{N}$. Based on the uniform mesh, we propose an explicit Euler-type method for (2.2) as follows:
@@
Let $c_{-1},c_{0},c_{1},c_{2},c_{3}>0$, $\kappa$, $\rho>1$. Given any initial value $Y_{0}=X_{0}>0$, the scheme (3.1) admits unique, positive numerical solutions $\{Y_{n}\}_{n\in\mathbb{N}}$ for any step-size $h=\tfrac{T}{N}>0$. In other words, the scheme (3.1) is positivity preserving unconditionally.
Proof: For $x>0$, we introduce the function
@@
Let $\kappa+1\geq 2\rho$ with $\kappa,\rho>1$. There exists a positive constant $L(\gamma)$, independent of the step-size $h>0$, such that
@@
where $\gamma\geq\tfrac{1}{2}$ is assumed to be arbitrarily large for the case $\kappa+1>2\rho$ and $\tfrac{1}{2}\leq\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ for the critical case $\kappa+1=2\rho$.
@@
There exist some positive constants $m_{1},m_{2}\leq 2\gamma+1$ such that
@@
where $\alpha$ is a positive constant satisfying $\alpha\geq\tfrac{1}{2}$. Here, the choice of $f_{h}$ and $g_{h}$ are nothing but a kind of taming strategy, which was firstly proposed in [11] and later extensively studied by many authors, see, e.g., [21, 3, 18, 19, 13, 14]. To validate Assumption 3.2, noticing that $2\kappa\alpha\geq\kappa>\rho$, simple calculations infer that for any $x\in(0,+\infty)$
@@
Thus, we have
$|f_{h}(x)|\leq c_{2}h^{-\frac{1}{2}}\quad\text{and}\quad|g_{h}(x)|\leq c_{3}h^{-\frac{\rho}{2\kappa}}\leq c_{3}T^{\frac{1}{2}-\frac{\rho}{2\kappa}}h^{-\frac{1}{2}},$ (3.11)
@@
$\big{|}g(x)-g_{h}(x)\big{|}$ $=\tfrac{c_{3}h^{\alpha}x^{{\color[rgb]{0,0,0}2\kappa\alpha+\rho}}}{1+h^{\alpha}x^{2\kappa\alpha}}\leq c_{3}h^{\alpha}x^{{\color[rgb]{0,0,0}2\kappa\alpha+\rho}}.$ (3.15)
Let $m_{1}=2(2\alpha+1)\kappa$ and $m_{2}=2({\color[rgb]{0,0,0}2\kappa\alpha+\rho})$. Clearly, $m_{2}\leq 4\alpha\kappa+\kappa+1\leq m_{1}$. Since $\alpha\geq\tfrac{1}{2}$, (3.8) holds true on an additional condition $12(2\alpha+1)\kappa\leq 2\gamma+$ 1.
@@
Let Assumptions 3.2 and 3.3 hold. Then for any $12\leq p\leq 2\gamma+$ 1 and for any step-size $h>0$ it holds that
@@
Next, we introduce a time-continuous version of $Z_{n+1}$ in each time interval $[t_{n},t_{n+1}]$ as follows:
@@
Note that $\widetilde{Z}_{t_{n}}^{(n)}=Y_{n}\neq Z_{n}=\widetilde{Z}_{t_{n}}^{(n-1)}$. Owing to (3.6), a rough moment bound can be derived for $\widetilde{Z}_{t}^{(n)}$. Indeed, by the Jensen inequality, for any $p\geq 1$, we have
@@
where $\varepsilon\geq 1+2c_{-1}h>1$ is a positive constant specified later and $\hat{C}$$,\tilde{C}_{\varepsilon}$ are positive constants independent of $h$. By noting $\varepsilon\geq 2c_{-1}h$, we infer
$\mathbb{E}\bigg{[}\Big{(}\varepsilon+\big{|}\widetilde{Z}_{t_{n+1}}^{(n)}\big{|}^{2}\Big{)}^{p}\bigg{]}$ $=\mathbb{E}\bigg{[}\Big{(}\varepsilon+\big{|}Y_{n+1}-c_{-1}hY_{n+1}^{-1}\big{|}^{2}\Big{)}^{p}\bigg{]}$
$=\mathbb{E}\bigg{[}\Big{(}\varepsilon+|Y_{n+1}|^{2}+c_{-1}^{2}h^{2}|Y_{n+1}|^{-2}-2c_{-1}h\Big{)}^{p}\bigg{]}$
$\geq\mathbb{E}\big{[}\big{|}Y_{n+1}\big{|}^{2p}\big{]}.$ (3.21)
Therefore, we deduce from (3.20) that
@@
By iteration, we get, for any $p\geq 1$ and $0\leq n\leq N$,
$\mathbb{E}\big{[}\big{|}Y_{n}\big{|}^{2p}\big{]}\leq\hat{C}^{n}\mathbb{E}\big{[}\big{|}Y_{0}\big{|}^{2p}\big{]}+\tilde{C}_{\varepsilon}\sum_{i=0}^{n-1}\hat{C}^{i}<\infty,$ (3.23)
which, together with (3.20), infers that, for any $p\geq 1$,
@@
We emphasize that the above rough moment bound, depending on $N$, is not proved to be uniformly bounded with respect to $N$. Nevertheless, it suffices to allow the use of the Gronwall lemma later to arrive at uniform moment bounds. To do this, we denote
$V_{\varepsilon,q}(x):=\big{(}\varepsilon+|x|^{2}\big{)}^{q},\quad q,x\in\mathbb{R},$ (3.25)
where we recall $\varepsilon\geq 1+2c_{-1}h$. It can be easily checked that for any $q,q^{\prime}\in\mathbb{R}$,
@@
$V_{\varepsilon,q}^{\prime\prime}(x)$ $=2q\big{(}\varepsilon+|x|^{2}\big{)}^{q-2}\big{[}\varepsilon+(2q-1)|x|^{2}\big{]}$ $\leq 2q\cdot\max{\{1,2q-1\}}\cdot V_{\varepsilon,q-1}(x)$
@@
Equipped with these properties, we can start the estimation of $\widetilde{Z}_{t}^{(n)}$. The Itô formula together with (3.28) and (3.29) yields that for any $q\geq 1$,
@@
where we also used (3.19) in the last step. In view of (3.7), for any $q\leq\gamma+\tfrac{1}{2}$, where $\gamma$ is assumed to be arbitrarily large for the case $\kappa+1>2\rho$ and $\tfrac{1}{2}\leq\gamma\leq\tfrac{c_{2}}{c_{3}^{2}}$ for the critical case $\kappa+1=2\rho$, one deduces that
@@
Recalling (3.19) and (3.25), one knows that the process $\widetilde{Z}_{s}^{(n)}$ is $\mathcal{F}_{s}$-measurable, $Y_{n}$ is $\mathcal{F}_{t_{n}}$-measurable by construction and both are integrable in view of (3.2) and (3.24). This ensures that the following expectation of the Itô stochastic integral vanishes:
$\mathbb{E}\Bigg{[}\int_{t_{n}}^{t}2q\widetilde{Z}_{s}^{(n)}V_{\varepsilon,q-1}(\widetilde{Z}_{s}^{(n)})g_{h}(Y_{n})\,\text{d}W_{s}\Bigg{]}=0.$ (3.35)
By observing that
$|Y_{n}|^{2q}\leq\big{(}\varepsilon+|Y_{n}|^{2}\big{)}^{q}=V_{\varepsilon,q}(Y_{n})=V_{\varepsilon,q}(\widetilde{Z}_{t_{n}}^{(n)}),$ (3.36)
we take expectations on both sides of (3.34) to show